A ten-foot-long board is cut into three pieces. The second piece is half as long as the first. The third piece is 4 1/2 feet longer than the second. How long is the first piece?

A. 1 1/10 feet
B. 2 1/5 feet
C. 2 3/4 feet
D. 3 1/3 feet

Answer by tutorcecilia(2144)

Piece #1 + Piece #2 + Piece #3=10 feet
.
Piece #1=x
Piece#2=1/2x=0.5x
Piece#3=.05x+ 4 1/2=.05x+4.5
Total board = 10
.
Piece #1 + Piece #2 + Piece #3=10 feet [plug-in the values and solve for x]
(x)+(0.5x)+(0.5x+4.5)=10
2x+4.5=10
2x+4.5-4.5=10-4.5
2x=5.5
2x/2=5.5/2
x=2.75 [plug- this back into the original equation]
.
(x)+(0.5x)+(0.5x+4.5)=10
(2.75)+(0.5)(2.75))+(0.5)(2.75)+4.5)=10
10=10 [checks out]